MICROWAVE SPECTROSCOPY AT THE DISSOCIATION LIMIT

\underline{ALAN CARRINGTON}, Department of Chemistry, University of Southampton, UK.

An ion beam technique has been developed which allows energy levels lying close to the dissociation limit to be studied by microwave spectroscopy. Electric field dissociation provides a high degree of state selectivity, with corresponding sensitivity. The technique was originally developed to study the spectra of H_{2}^{(+)} and its isotopes, but has since been extended to a number of rare gas dimer ions, and one triatomic ion. All of the species studied have open-shell ground states, and exhibit different angular momenta coupling schemes as dissociation is approached. The details will be discussed in this talk.